# On the Sequence Evaluation based on Stochastic Processes

### v3

@@ 2.1 Stochastic representation
$\widehat{\Sigma}=\Big{(}\sum_{i=1}^{n}(T_{i}-1)\Big{)}^{-1}\Big{(}\sum_{i=1}^{n}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top}\Big{)}.$
The training of the SP Encoder and the definition of the SPM are thus based on the MLE of $\Sigma$.
2.2 Encoders
The encoder architecture consists of two components: a frozen, pretrained GPT-2 model from the Hugging Face library and a trainable multi-layer perceptron (MLP) network. We extract the hidden state corresponding to the end-of-sentence (EOS) token from the last layer of the GPT-2 model. This hidden state serves as the input to a four-layer MLP, which is trained to map the input into the latent space. The purpose of the encoder is to learn a nonlinear mapping from the raw input space to the latent space, denoted as $f_{\theta}:{\mathcal{X}}\to{\mathcal{S}}$. In practice, we use two loss functions to train the encoder: the contrastive learning loss $L_{\rm CL}$ and the negative log-likelihood loss $L_{\rm NLL}$. The contrastive learning loss $L_{\rm CL}$ is primarily designed to enhance the encoder’s ability to distinguish between positive and negative samples from the data, following the approach in (van den Oord et al., 2018). In contrast, the negative log-likelihood loss $L_{\rm NLL}$ directly enforces the latent space representations to follow a specified distribution according to the likelihood function. While Wang et al. (2022) utilized contrastive learning for the encoder, our SP Encoder employs the negative log-likelihood loss, which offers certain advantages over their approach, both theoretically and empirically. For comparison purposes, we first introduce the contrastive learning encoder as presented in Wang et al. (2022).
2.2.1 Contrastive Learning Encoder (CL Encoder)
In the context of Wang et al. (2022), a structure assumption $\Sigma={\mathbf{I}}_{d}$ was imposed, where ${\mathbf{I}}_{d}$ is the $d$-dimensional identity matrix. For arbitrary starting point ${\mathbf{s}}_{0}$ at time $t=0$ and ending point ${\mathbf{s}}_{T}$ at time $t=T$, the marginal distribution of ${\mathbf{s}}_{t}$ at time $t$ is given by ${\mathbf{s}}_{t}\mid{\mathbf{s}}_{0},{\mathbf{s}}_{T}\sim N((1-t/T){\mathbf{s}}_{0}+(t/T){\mathbf{s}}_{T},[t(T-t)/T]{\mathbf{I}}_{d})$. Consider any triplet of observations $({\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3})$ with ${\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3}\in{\mathcal{X}}$. The goal is to ensure that $f_{\theta}({\mathbf{x}}_{2})$ follows the above marginal distribution with starting point $f_{\theta}({\mathbf{x}}_{1})$ and ending point $f_{\theta}({\mathbf{x}}_{3})$. For a sequence of observations $({\mathbf{x}}_{0},\ldots,{\mathbf{x}}_{T})$, let $B=\{({\mathbf{x}}_{0},{\mathbf{x}}_{t},{\mathbf{x}}_{T})\}$ be a batch consisting of randomly sampled positive triplets $({\mathbf{x}}_{0},{\mathbf{x}}_{t},{\mathbf{x}}_{T})$ with $0<t<T$. Then, the contrastive learning loss function is defined as
$L_{\rm CL}=\mathrm{E}\left[-\log\frac{\exp(d({\mathbf{x}}_{0},{\mathbf{x}}_{t},{\mathbf{x}}_{T};f_{\theta}))}{\sum_{({\mathbf{x}}_{0},{\mathbf{x}}_{t^{\prime}},{\mathbf{x}}_{T})\in B}\exp(d({\mathbf{x}}_{0},{\mathbf{x}}_{t^{\prime}},{\mathbf{x}}_{T};f_{\theta}))}\right],$
@@ 2.2.1 Contrastive Learning Encoder (CL Encoder)
$d({\mathbf{x}}_{0},{\mathbf{x}}_{t},{\mathbf{x}}_{T};f_{\theta})=-(2\sigma^{2})^{-1}\lVert f_{\theta}({\mathbf{x}}_{t})-(1-t/T)f_{\theta}({\mathbf{x}}_{0})-(t/T)f_{\theta}({\mathbf{x}}_{T})\rVert_{2}^{2},~{}~{}\sigma^{2}=t(T-t)/T.$
However, there are two theoretical drawbacks to this contrastive learning approach. First, the assumption $\Sigma={\mathbf{I}}_{d}$ implies independence and homogeneity among the dimensions of the encoded sequence, which may not capture the true structural dependencies present in the sequence. Second, using the marginal distribution neglects the time-dependent structure of the stochastic process, limiting the model’s ability to capture temporal dependencies.
Empirically, it has been observed that only one dimension of the encoder’s output is effective, which is undesirable for representing complex sequences. To address these issues and incorporate both structural and temporal dependencies, we propose our novel SP Encoder, which minimizes the negative log-likelihood loss. This approach allows us to model the latent space more accurately by considering the full covariance structure and the temporal dynamics of the data, leading to improved theoretical foundations and empirical performance over the contrastive learning method.
2.2.2 Stochastic Process Encoder (SP Encoder)
@@
Consider a multi-domain problem with $m$ domains ${\mathcal{X}}_{1},{\mathcal{X}}_{2},\ldots,{\mathcal{X}}_{m}$ each associated with domain-specific true structural parameters $\Sigma_{1},\Sigma_{2},\ldots,\Sigma_{m}$, respectively. For each domain ${\mathcal{X}}_{j}$, we have $n_{j}$ independent raw inputs ${\mathbf{x}}_{j1},\ldots,{\mathbf{x}}_{jn_{j}}$. We define the encoded sequences as $\bar{\mathbf{s}}_{ji}^{\theta}=f_{\theta}({\mathbf{x}}_{ji})$ for $j=1,\ldots,m$ and $i=1,\ldots,n_{j}$, where $f_{\theta}$ is the encoder parameterized by $\theta$. When the encoder parameters reach their optimal values $\theta^{*}$, the sequences $[\bar{{\mathbf{s}}}_{ji}^{\theta^{*}}]_{i=1}^{n_{j}}$ are expected to be i.i.d. samples from BBs with parameters $\Sigma_{j}$ for each domain ${\mathcal{X}}_{j}$.
@@
We employ the negative log-likelihood (NLL) as the loss function to train the encoder. According to Proposition 1, for each $\theta$, the negative log-likelihood for domain ${\mathcal{X}}_{j}$ depends on $\Sigma_{j}$ and the inputs $[{\mathbf{x}}_{ji}]_{i=1}^{n_{j}}$ through the expression $\sum_{i=1}^{n_{j}}(T_{i}-1)\log(\lvert\Sigma_{j}\rvert)+\sum_{i=1}^{n_{j}}\mathrm{tr}(\Sigma_{j}^{-1}({\mathbf{s}}_{i}^{\theta}-\bm{\mu}_{i}^{\theta})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}^{\theta}-\bm{\mu}_{i}^{\theta})^{\top})$. We consider the following training process.
@@
Batch Processing: We divide the inputs $[{\mathbf{x}}_{ji}]_{i=1}^{n_{j}}$ into several batches. For each batch ${\mathcal{B}}$, we compute the batch loss using the current estimate $\widehat{\Sigma}_{j}$ of $\Sigma_{j}$: $\sum_{i\in{\mathcal{B}}}\mathrm{tr}(\widehat{\Sigma}_{j}^{-1}({\mathbf{s}}_{i}^{\theta}-\bm{\mu}_{i}^{\theta})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}^{\theta}-\bm{\mu}_{i}^{\theta})^{\top})$. This loss function measures how well the encoded sequences fit the assumed BB model with the current structural parameter estimate.
@@
Handling Large Sequences: When the sequence lengths $T_{i}$ are large, computing the full loss can be computationally intensive. To address this, we randomly sample a triplet of time points $t=(t_{1},t_{2},t_{3})$ with $11\leq t_{1}<t_{2}<t_{3}\leq T_{i}-$ 1. We extract the corresponding sub-matrices $[{\mathbf{s}}_{i}^{\theta}]_{t}$ and $[\bm{\mu}_{i}^{\theta}]_{t}$ of size $d\times 3$ from ${\mathbf{s}}_{i}^{\theta}$ and $\bm{\mu}_{i}^{\theta}$, respectively. Let $[\Sigma_{T_{i}}]_{t}$ be the $3\times 3$ sub-matrix of $\Sigma_{T_{i}}$ corresponding to the selected time points. The loss for each $i$ in the batch becomes $\mathrm{tr}(\widehat{\Sigma}_{j}^{-1}([{\mathbf{s}}_{i}^{\theta}]_{t}-[\bm{\mu}_{i}^{\theta}]_{t})[\Sigma_{T_{i}}]_{t}^{-1}([{\mathbf{s}}_{i}^{\theta}]_{t}-[\bm{\mu}_{i}^{\theta}]_{t})^{\top})$. This approach reduces computational complexity while still capturing temporal dependencies at selected time points.
@@
Updating Structural Parameters: After processing all batches for ${\mathcal{X}}_{j}$, we update the estimate of $\Sigma_{j}$ using the MLE: $\widehat{\Sigma}_{j}=[\sum_{i=1}^{n_{j}}(T_{i}-1)]^{-1}[\sum_{i=1}^{n_{j}}({\mathbf{s}}_{i}^{\theta}-\bm{\mu}_{i}^{\theta})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}^{\theta}-\bm{\mu}_{i}^{\theta})^{\top}]$. This update aggregates information from all sequences in the domain to refine the structural parameter estimate.

### v4

@@ 3.3 Alignment in latent space
$\begin{aligned} \widehat{\Sigma}=\Big(\sum_{i=1}^{n}(T_{i}-1)\Big)^{-1}\Big(\sum_{i=1}^{n}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\boldsymbol{\mu}_{i})^{\top}\Big).\end{aligned}$
The definition of the BBScoreV2 is therefore derived from the MLE of $\Sigma$. Consider the sequence $\bar{{\mathbf{s}}}=(s_{0},\ldots,s_{T})$, with $\mathbf{s}$} and $\boldsymbol{\mu}$ defined as before. To evaluate the coherence of the sequence from a domain with unknown parameters $\Sigma$, a natural approach is to compute its density under the assumed model. If $\bar{{\mathbf{s}}}$ is a BB with covariance $\Sigma$, then by Proposition 1, the log density of $\bar{{\mathbf{s}}}$ is given by Equation 4. To remove the length sensitive term in the log density, we design a standardized score for practical purposes, and define the score as following:
Definition (BBScoreV2).
Let $\widehat{\Sigma}$ be the estimate of $\Sigma$ from Proposition 2. The metric BBScoreV2 is defined as
${\mathcal{B}}^{+}(\bar{{\mathbf{s}}}|\widehat{\Sigma})=\log p(\bar{{\mathbf{s}}}|\Sigma)/[d(T-1)].$
Given an accurate estimate $\widehat{\Sigma}$ of the true covariance $\Sigma$, and assuming the input sequence $\bar{{\mathbf{s}}}$ originates from a BB process with covariance $\Sigma$, a lower BBScoreV2 value signifies a decreased likelihood of $\bar{{\mathbf{s}}}$ being generated under $\Sigma$. Conversely, if the representation encodes a better temporal and structural information, such as encoded from a more coherent article, the probability density will be higher, resulting in a larger BBScoreV2.
In a summary, BBScoreV2 is novel in two key aspects. First, by utilizing the temporal covariance matrix $\Sigma_{T}$, the BBScoreV2 captures the time-dependent structure inherent in the sequence, which is essential for accurately assessing sequence temporal property, such as coherence. Second, the inclusion of the covariance matrix $\Sigma$ allows the BBScoreV2 to account for structural dependencies among the latent dimensions, providing a more comprehensive evaluation of the sequence’s adherence to the assumed stochastic process.
4 Experiments and Problems
To understand the spatial and temporal information encoded in stochastic representations, we experimentally designed latent space visualization experiments. Subsequently, we evaluate BBScoreV2 to demonstrate its utility in downstream tasks that leverage this encoded information. Our experiments are designed to address the following three key research questions (Q):
  • Q1: How is stochastic representation learning achieved, and what makes it effective? In Section 5.1, we analyze the spatial structure of the latent space.
  • Q2: Can BBScoreV2 capture correct temporal information and assess document coherence? In Section 5.2, we examine its performance on standard shuffle tasks (indicative of temporal understanding) and also comparing the coherence of articles of varying lengths—an evaluation that current state-of-the-art methods often cannot perform effectively.
  • Q3: Can we use BBScoreV2 to detect AI-generated text from human-written ones ? In Section 5.3, we explore whether BBScoreV2 can effectively distinguish between human-written text and text generated by AI. We also compare its performance with other baselines.
To validate the above question, we design the following experiments. Moreover, in Section D, we describes the dataset utilized in these experiments and how we construct the input.
Global discrimination.
We employed the Shuffle Test (Barzilay and Lapata, 2005; Moon et al., 2019) to assess BBScoreV2’s ability to evaluate temporal information and discriminate global coherence. It involves randomly permuting sentences within a document to create an incoherent version, which is then compared against the original. Specifically, for each article, we generated 20 unique shuffled copies by permuting entire sentence blocks of varying sizes (1, 2, 5, and 10 sentences).
Mixed Shuffled test.
Building upon the standard Shuffle Test, we introduced a more challenging variant called the Mixed Shuffle Test. In this setup, BBScoreV2 of an original (unshuffled) article is compared against BBScoreV2 of shuffled articles drawn from the entire dataset, rather than solely against its own shuffled versions. A robust and general-purpose scoring mechanism should consistently identify the original, unshuffled article as more coherent in these broader comparisons.
Human-AI text discrimination.
We leverage the HC3 Q&A dataset (Guo et al., 2023) to train the encoder exclusively on human-generated answers, and subsequently apply it to unseen Q&A pairs generated by both humans and ChatGPT. After deriving the stochastic representations, we compute the BBScoreV2 for each Q&A pair. We evaluate multiple encoder backbones to examine the impact of the raw embeddings. Additionally, we train an encoder on the WikiSection dataset and evaluate it using the Wikipedia subset of HC3. Experiments are conducted under both the full Q&A and answer-only settings to determine if the BBScoreV2 can effectively discriminate between ChatGPT-generated and human-written texts.
5 Results
5.1 Latent space structure analysis
Figure 2: PCA analysis of raw LM embeddings and CL encoder representations. (A.1) Projection of raw LM embeddings for an unshuffled article onto the first two PCs. The color gradient, from light to dark red, indicates the sequential position of each token within the article. (A.2, A.3) raw LM embeddings for shuffled versions of the same article. (B) Mean PC1 and PC2 values for all articles are plotted, with each article represented by a dot. (C.1) Latent representations from the CL encoder for an unshuffled article, where the color gradient (light to dark blue) signifies the component’s position in the article sequence. (C.2, C.3) Latent patterns observed in shuffled versions of the same article. (D) Visualization of the latent trajectories for all articles.
Theoretically, transformer-based LLMs are argued to map articles to a latent representation that tends to form clusters. The structural properties of these clusters are believed to reflect underlying similarities and properties present in the original articles Geshkovski et al. (2023).
Experimentally, we also find such clustered property. We first visualized the raw embeddings of each article from the frozen GPT-2 model. In Figure 2 (A.1), these embeddings are projected onto their joint first two principal components (PCs), derived using PCA computed from the latents of unshuffled articles. The color gradient, from light to dark red, represents the token’s sequential position within the article, from beginning to end. Notably, as shown in (A.2) and (A.3) for shuffled versions of an article, the distinct clustering property persists. This persistence, despite the disruption of sequential order, suggests that these raw LM embeddings do not clearly and inherently encode temporal information. To further substantiate this, Figure 2 (B) plots the mean values of the first two PCs for the embeddings of each article, illustrating a tendency for articles from the same dataset to cluster together based on their raw LM embeddings.
Subsequently, to determine the information learned by the MLP layers in our CL encoder, we analyzed its outputted stochastic representations. Figure 2 (C.1) displays these MLP-processed latents projected via PCA. Here, a color gradient from light to dark blue indicates the component’s position within the article’s sequence (from begin to the end). This visualization reveals a clear temporal progression in the latent space for the original, unshuffled article. In stark contrast, Figures (C.2) and (C.3), which depict shuffled versions of the same article, demonstrate that the CL encoder’s representations clearly reflect this violation of temporal order; the clear sequential pattern observed in (C.1) is visibly disrupted. Furthermore, Figure 2 (D) presents the projection of latent trajectories for all articles. This visualization further validates our assertion that the CL encoder effectively learns and represents temporal sequence information, unlike the raw LLM embeddings.
Based on these findings, we show that the CL encoder effectively encodes temporal information into the representation. Furthermore, by evaluating the temporal structure, we can infer properties of the original articles—such as coherence—which are quantified by BBScore+ and will be systematically discussed in the following sections.
5.2 Article coherence evaluation
Methods Acc. (Shuffle Task) Acc. (Mixed Shuffle Task)
Entity Grid (Barzilay and Lapata, 2005) 85.73 82.79 75.81 64.65 46.10 52.29 53.69 63.02
Unified Coherence (Moon et al., 2019) 99.73 97.86 96.90 96.09
BBScore (Sheng et al., 2024) 83.39 80.71 79.36 78.66 22.37 24.94 23.84 19.69
BBScoreV2 (GPT2-124M) 99.03 98.11 98.02 98.17 94.78 89.24 79.64 70.83
BBScoreV2 (LLaMA3-1B) 99.16 98.37 97.99 97.87 94.53 87.86 76.95 71.13
BBScoreV2 (LLaMA3-3B) 99.57 98.74 98.14 98.74 94.97 86.34 73.88 68.87
Table 1: Results of Global shuffle tasks on WikiSection. $\mathcal{D}_{b=i}$, $i=1,2,5,10$ refers to datasets constructed with varying levels of block shuffling.
As shown in Tables 1, we first implement global discrimination tasks on WikiSection. In this task, BBScoreV2 significantly outperforms the BBScore and SOTA results. (See Appendix E for more details on methods we compared to.) The SOTA method, developed using a complex network structure and trained on unshuffle-shuffle data pairs, serves as a robust baseline. Our results demonstrate that BBScoreV2 surpasses the SOTA method in global discrimination tasks with larger block sizes, underscoring its potential to capture more globalized temporal properties.
In shuffle tasks, most current high-performance methods, including the SOTA approach, rely on pairwise training and are unable to effectively compare articles of different lengths, as these models are typically constructed based on sentence-wise matching and comparisons. However, in the Mixed Shuffle test which evaluate the metric robustness across different articles, as shown in Table 1, BBScoreV2 surpasses these SOTA method by generating a metric that can be compared across different articles. We use the basic entity-grid method (Barzilay and Lapata, 2005) as a baseline and the result highlights that our score enables article-wise comparison. It also demonstrates significant potential in more complex tasks. Additionally, BBScoreV2 outperforms the BBScore in this article-wise comparison, underscoring a key contribution of our design—mitigating the effect of article length on score evaluation. This property allows for a more general comparison across diverse articles.
We also explore the effect of different LLM backbones. We tested our model using LLaMA3-1B and LLaMA3-3B, with GPT2-124M which is the LLM model used in the main section. As summarized in Table 1, we find: 1) In global shuffle task, LLaMA3-3B outperforms both GPT2-124M and the SOTA method, demonstrating its effectiveness in capturing global sequence structure; 2) In Mixed Shuffle Task, LLaMA3-3B surpasses GPT2-124M for smaller blocks (b=1), but its performance decreases for larger blocks (b=2, b=5, b=10). This suggests a trade-off where larger models excel at capturing local details (b=1) but might sacrifice robustness for global structures (b=10). This insight highlights an intriguing direction for future exploration — different LMs may facilitate learning stochastic representations in task-specific ways.
Moreover, we evaluate the robustness of the encoded stochastic representation on a broader dataset. As shown in Table 2, we train the encoder on WikiText and evaluate it on WikiSection (see Appendix D for details and comparisons about the datasets). The results indicate that our method remains highly robust in this O.O.D. setting, suggesting that the structural and temporal information captured by our model reflects fundamental patterns that generalize across different datasets.
Methods Shuffle Test tasks (O.O.D.)
$\mathcal{D}_{b=1}$ $\mathcal{D}_{b=2}$ $\mathcal{D}_{b=5}$ $\mathcal{D}_{b=10}$
Unified Coherence 60.02 9.63 44.80 66.51
BBScore 70.32 72.09 76.84 77.73
BBScoreV2 91.30 87.22 86.14 88.18
Table 2: O.O.D. Task. Encoder was trained on the WikiText and evaluated on Shuffle Test tasks using the same WikiSection data to assess their performance.
5.3 Human-AI discrimination tasks
In this task, we hypothesize that human writing, compared to AI-generated text, displays temporal dynamics and structural patterns similar to those observed in other human-written articles. Specifically, we propose that an encoder trained on a human-written dataset will more accurately capture the characteristics of human writing than those of AI-generated text, resulting in a higher likelihood for human-authored content. As shown in Figure 3, BBScoreV2 consistently outperforms BBScore across all experimental settings. Notably, GPT2 (124M) surpasses the larger backbone models, suggesting that the quality of the learned stochastic representation does not necessarily improve with increased model size.Instead, it is the MLP module that plays the central role in shaping the stochastic representation. Among the models evaluated, GPT2 features a smaller hidden dimension of $768$, whereas both LLaMA3 and Qwen3 utilize larger hidden dimensions of $2048$. This variation suggests that the hidden dimension of the backbone model may influence performance on this task.
Figure 3: Compare different LM backbones.
Next, we use a WikiSection trained encoder to detect ChatGPT-generated answer in the Wikipedia subset of HC3. The results are shown in Table 3 under both Q&A and answer-only settings. To further highlight the flexibility and competitive performance of BBScoreV2 compared to LLM-based models, we assessed the perturbation discrepancy metric proposed in DetectGPT (Mitchell et al., 2023), which has high performance in AI detection tasks. Our results reveal that BBScoreV2 surpasses DetectGPT when using a comparable number of model inferences. DetectGPT’s performance is influenced by a hyperparameter—the number of perturbations—which directly affects both the number of model inferences and the computational complexity. As shown in Table 7 in the Appendix, we tested cases with 1 and 10 perturbations. With 1 perturbation, DetectGPT’s accuracy was approximately 64%, lower than BBScoreV2’s 70%, while requiring twice the number of model inferences per text. With 10 perturbations, DetectGPT’s accuracy increased to 84%, but this required 11 model inferences per text, making it significantly more computationally intensive than BBScoreV2.
Methods HC3 (w/o Q&A) HC3 (w/ Q&A)
BBScore 37.53 31.47
DetectGPT 64.30 63.30
BBScoreV2 70.67 69.71
Table 3: Accuracy of the Human-AI discrimination task.
5.4 Ablation analysis on CL encoder
As previously discussed, the CL encoder relies on a critical assumption of the independence and homogeneity among the dimensions of the encoded sequence which is $\Sigma={\mathbf{I}}_{d}$. To further examine this assumption, we employ two alternative methods:
1) A likelihood-based encoder, SP Encoder (see Appendix A) whose loss function is defined based on the likelihood of the Brownian bridge:
[…]
$+\sum_{j=1}^{m}\sum_{i=1}^{n_{j}}\mathrm{tr}(\Sigma_{j}^{-1}({\mathbf{s}}_{i}^{\theta}-\boldsymbol{\mu}_{i}^{\theta})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}^{\theta}-\boldsymbol{\mu}_{i}^{\theta})^{\top}).$
2) A contrastive loss-based encoder, whose loss function is AnInfoNCE Rusak et al. (2024) which is capable of learning $\Sigma$ during training (see Appendix B for details). The loss function $L_{\rm AnInfoNCE}$ is defined as following:
$\mathrm{E}\left[-\log\frac{\exp(d^{*}({\mathbf{x}}_{0},{\mathbf{x}}_{t},{\mathbf{x}}_{T};f_{\theta}))}{\sum_{({\mathbf{x}}_{0},{\mathbf{x}}_{t^{\prime}},{\mathbf{x}}_{T})\in B}\exp(d^{*}({\mathbf{x}}_{0},{\mathbf{x}}_{t^{\prime}},{\mathbf{x}}_{T};f_{\theta}))}\right],$
@@ 5.4 Ablation analysis on CL encoder
$d^{*}({\mathbf{x}}_{0},{\mathbf{x}}_{t},{\mathbf{x}}_{T};f_{\theta})$
$=-\frac{\lVert f_{\theta}({\mathbf{x}}_{t})-\frac{T-t}{T}f_{\theta}({\mathbf{x}}_{0})-\frac{t}{T}f_{\theta}({\mathbf{x}}_{T})\rVert_{\hat{\Lambda}}^{2}}{2t(T-t)/T}.$
and $\hat{\Lambda}$ is a trainable diagonal scaling matrix. Let $\mathbf{v}=f_{\theta}({\mathbf{x}}_{t})-\frac{T-t}{T}f_{\theta}({\mathbf{x}}_{0})-\frac{t}{T}f_{\theta}({\mathbf{x}}_{T})$, then its corresponding norm is defined as:
$\lVert\mathbf{v}\rVert_{\hat{\Lambda}}^{2}=\mathbf{v}^{\mathrm{T}}\cdot\hat{\Lambda}\cdot\mathbf{v}$
As shown in Table 4, neither the likelihood-based nor the CL encoder with AnInfoNCE loss yields significant improvements in the shuffle test. This suggests that the MLP layers do not capture meaningful structural correlations across latent dimensions—a phenomenon also noted by Wang et al. (2022)—and instead primarily reconstruct temporal information, which validate our assumptions on CL encoder training. The lack of performance improvement may also suggest that the pertinent correlation structure is likely inherent either within the statistical properties of the article domain or already captured within the high-dimensional embedding space of the pre-trained language model as we seen in the cluster analysis in Fig 2.
Loss Type $\mathcal{D}_{b=1}$ $\mathcal{D}_{b=2}$ $\mathcal{D}_{b=5}$ $\mathcal{D}_{b=10}$
AnInfoNCE 94.63 91.05 92.13 91.70
Likelihood 94.42 92.90 90.77 86.69
Ours 99.03 98.11 98.02 98.17
Table 4: Comparison of Model Performance with different loss function on WikiSection Dataset.
5.5 Computation efficiency analysis
We specifically analyze the computation efficiency of BBScoreV2, as shown in Figure 4, the y-axis represents computation time, while the x-axis indicates article length. The theoretical computational complexity of BBScoreV2 is $O(T^{2})$, primarily due to matrix multiplications inherent in its definition. This complexity is fundamental to fully leveraging temporal information for sequence evaluation. Empirically, the observed computation time is slightly better than the theoretical prediction, thanks to the computational acceleration. These results demonstrate that BBScoreV2 is not only feasible for real-time applications but also retains its robust evaluation capabilities.
Figure 4: The computation time of BBScoreV2 for different article lengths. It reveals a quadratic relationship (experimentally 1.57, theoretically 2) between article length and computation time, with each article processed in approximately $\sim 10^{-3}$ seconds.
6 Conclusion
In this paper, we present both a theoretical and empirical investigation into the structural and temporal properties encoded in stochastic representations of latent trajectories for NLP tasks. We analyze and visualize these properties, and introduce BBScoreV2—a novel, length-invariant metric designed to quantify such information. First we present the learned representations recovers the time dependency of the input sequences. Then validated through shuffled and mixed-shuffle tests, we show that BBScoreV2 exhibits strong performance in capturing temporal structure and generalizes effectively to out-of-distribution tasks, suggesting that these properties reflect domain-independent textual signals. Moreover, BBScoreV2 shows promising capability in distinguishing human-written from AI-generated text by leveraging encoded structural and temporal features.
Looking ahead, we aim to extend BBScoreV2 to multi-domain tasks such as domain identification, and to exploit its length-insensitive nature to develop generative models that maintain semantic coherence across varying sequence lengths. Its computational efficiency (see Fig. 4) also makes it suitable for large-scale applications. Finally, inspired by Albergo et al. (2023); Albergo and Vanden-Eijnden (2023), we plan to explore more expressive bridge processes to further enhance the representational capacity of the latent space and enable richer downstream analysis and generation.
7 Limitations
First, our current study is constrained by limited computational resources and the lack of human-annotated data, which prevents us from evaluating BBScoreV2 against human preference—a key limitation in assessing its alignment with human judgment. Additionally, in the Human-AI text discrimination task, we were unable to evaluate the metric on a broader range of datasets or conduct more extensive comparisons across more baselines. These limitations suggest directions for future work involving large-scale human evaluation and broader benchmarking.
Appendix A Appendix: SP Encoder
A.1 Definition
@@
Consider a multi-domain problem with $m$ domains ${\mathcal{X}}_{1},{\mathcal{X}}_{2},\ldots,{\mathcal{X}}_{m}$ each associated with domain-specific true structural parameters $\Sigma_{1},\Sigma_{2},\ldots,\Sigma_{m}$, respectively. For each domain ${\mathcal{X}}_{j}$, we have $n_{j}$ independent raw inputs ${\mathbf{x}}_{j1},\ldots,{\mathbf{x}}_{jn_{j}}$. We define the encoded sequences as $\bar{\mathbf{s}}_{ji}^{\theta}=f_{\theta}({\mathbf{x}}_{ji})$ for $j=1,\ldots,m$ and $i=1,\ldots,n_{j}$, where $f_{\theta}$ is the encoder parameterized by $\theta$. When the encoder parameters reach their optimal values $\theta^{*}$, the sequences $\bar{{\mathbf{s}}}_{ji}^{\theta^{*}}]_{i=1}^{n_{j}}$ are expected to be i.i.d. samples from BBs with parameters $\Sigma_{j}$ for each domain ${\mathcal{X}}_{j}$.
@@
We employ the negative log-likelihood (NLL) as the loss function to train the encoder. According to Proposition 1, for each $\theta$, the negative log-likelihood for domain ${\mathcal{X}}_{j}$ depends on $\Sigma_{j}$ and the inputs ${\mathbf{x}}_{ji}]_{i=1}^{n_{j}}$ through the expression $\sum_{i=1}^{n_{j}}(T_{i}-1)\log(\lvert\Sigma_{j}\rvert)+\sum_{i=1}^{n_{j}}\mathrm{tr}(\Sigma_{j}^{-1}({\mathbf{s}}_{i}^{\theta}-\boldsymbol{\mu}_{i}^{\theta})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}^{\theta}-\boldsymbol{\mu}_{i}^{\theta})^{\top})$. We consider the following training process.
@@
Batch Processing: We divide the inputs ${\mathbf{x}}_{ji}]_{i=1}^{n_{j}}$ into several batches. For each batch $\mathcal{B}$}, we compute the batch loss using the current estimate $\widehat{\Sigma}_{j}$ of $\Sigma_{j}$: $\sum_{i\in{\mathcal{B}}}\mathrm{tr}(\widehat{\Sigma}_{j}^{-1}({\mathbf{s}}_{i}^{\theta}-\boldsymbol{\mu}_{i}^{\theta})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}^{\theta}-\boldsymbol{\mu}_{i}^{\theta})^{\top})$. This loss function measures how well the encoded sequences fit the assumed BB model with the current structural parameter estimate.
@@
Handling Large Sequences: When the sequence lengths $T_{i}$ are large, computing the full loss can be computationally intensive. To address this, we randomly sample a triplet of time points $t=(t_{1},t_{2},t_{3})$ with $1\leq t_{1}<t_{2}<t_{3}\leq T_{i}-1$. We extract the corresponding sub-matrices ${\mathbf{s}}_{i}^{\theta}]_{t}$ and $\boldsymbol{\mu}_{i}^{\theta}]_{t}$ of size $d\times 3$ from ${\mathbf{s}}_{i}^{\theta}$ and $\boldsymbol{\mu}_{i}^{\theta}$, respectively. Let $\Sigma_{T_{i}}]_{t}$ be the $3\times 3$ sub-matrix of $\Sigma_{T_{i}}$ corresponding to the selected time points. The loss for each $i$ in the batch becomes $\mathrm{tr}(\widehat{\Sigma}_{j}^{-1}([{\mathbf{s}}_{i}^{\theta}]_{t}-[\boldsymbol{\mu}_{i}^{\theta}]_{t})[\Sigma_{T_{i}}]_{t}^{-1}([{\mathbf{s}}_{i}^{\theta}]_{t}-[\boldsymbol{\mu}_{i}^{\theta}]_{t})^{\top})$. This approach reduces computational complexity while still capturing temporal dependencies at selected time points.
@@
Updating Structural Parameters: After processing all batches for ${\mathcal{X}}_{j}$, we update the estimate of $\Sigma_{j}$ using the MLE: $\widehat{\Sigma}_{j}=[\sum_{i=1}^{n_{j}}(T_{i}-1)]^{-1}[\sum_{i=1}^{n_{j}}({\mathbf{s}}_{i}^{\theta}-\boldsymbol{\mu}_{i}^{\theta})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}^{\theta}-\boldsymbol{\mu}_{i}^{\theta})^{\top}]$. This update aggregates information from all sequences in the domain to refine the structural parameter estimate.